How do you get the percentage of 85/110?
Im multiplying it by 100 but it comes out to this big decimal,
the answer is 77.27 but my question is how do you even get that number?
Or what are the steps exactly to getting that answer?

Found 2 solutions by jim_thompson5910, rfer:
Answer by jim_thompson5910(35256)

(85/110)*100 = (0.77272727272728)*100 = 77.272727272728 %

which rounds to 77.27%

Answer by rfer(16322)

85/110=0.7727=77.27%